Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 12:32 Saturday 17 April 2021 (16502) Print this report
Comment to Tightening the locating clamps for the piano wire. (16348)

With Hirata-san.

This is a report for the work carried out yesterday.

See pictures in album PR3 Remedying Work.

  • We locked the IRM and moved the three vertical OSEMs either into range or to a better position.
  • We released the IRM and measured a few transfer functions, which showed the payload is heathly.
  • For the purposes of checking reproducilility we measured the pitch of the mirror again using the good laser level; we got -3.3mrad, which is the value we were expecting.
Search Help